### v1

@@ 4 The magnetic field integral equation
where $\boldsymbol{u}:=\gamma_{N}^{+}\boldsymbol{e}$ is the unknown exterior Neumann trace of the scattered electric field, and $\boldsymbol{f}\in\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$ is given via the Dirichlet boundary condition (4) as
@@
Then, the variational formulation of (9) reads as: find $\boldsymbol{u}\in\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$ such that, for all $\boldsymbol{v}\in\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$
(10) $b(\boldsymbol{u},\boldsymbol{v})=\left\langle{\boldsymbol{f},\overline{\boldsymbol{v}}}\right\rangle_{\times,\Gamma},$
where the sesquilinear form $b:\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)\times\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)\to\mathbb{C}$ is defined by
$b(\boldsymbol{u},\boldsymbol{v}):=\left\langle{\left({\dfrac{1}{2}Id+C_{\kappa}}\right)\boldsymbol{u},\overline{\boldsymbol{v}}}\right\rangle_{\times,\Gamma}.$
On smooth surfaces, it is well-known that the operator $C_{\kappa}$ is compact, see, e.g., [29, Section 5.4] or [7, Lemma 11]. As a result, the self-duality of the space $\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$ ensures the coercivity of the sesquilinear form $b(\cdot,\cdot)$. More precisely, the following inf-sup condition is satisfied for all $\boldsymbol{u}\in\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$
(11) $\sup_{\boldsymbol{v}\in\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)\setminus\left\{{\boldsymbol{0}}\right\}}\dfrac{\left|{d(\boldsymbol{u},\boldsymbol{v})}\right|}{\left\lVert{\boldsymbol{v}}\right\rVert_{\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)}}\geq C_{D}\left\lVert{\boldsymbol{u}}\right\rVert_{\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)},$
where $d:\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)\times\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)\to\mathbb{C}$ is the sesquilinear form induced by the duality pairing ,×$\left\langle{\cdot,\cdot}\right\rangle_{\times,\Gamma}$, i.e.,
(12) $d(\boldsymbol{u},\boldsymbol{v}):=\left\langle{\boldsymbol{u},\overline{\boldsymbol{v}}}\right\rangle_{\times,\Gamma}.$
On polyhedral boundaries, however, the compactness of $C_{\kappa}$ does not necessarily hold true. In this paper, we exploit the compactness of the operator $\delta C_{\kappa}$ (cf. Lemma 3.2) and the $S_{i\kappa}$-coercivity stated in (8) to prove the coercivity of the sesquilinear form $b(\cdot,\cdot)$. For that reason, we split $b(\cdot,\cdot)$ into a principal part $a(\cdot,\cdot)$ and a compact sesquilinear form $t(\cdot,\cdot)$ as follows
@@
$a(\boldsymbol{u},\boldsymbol{v}):=\left\langle{\left({\dfrac{1}{2}Id+C_{i\kappa}}\right)\boldsymbol{u},\overline{\boldsymbol{v}}}\right\rangle_{\times,\Gamma},\qquad\quad t(\boldsymbol{u},\boldsymbol{v}):=\left\langle{\delta C_{\kappa}\boldsymbol{u},\overline{\boldsymbol{v}}}\right\rangle_{\times,\Gamma}.$
The $S_{i\kappa}$-coercivity of the operator $\tfrac{1}{2}Id+C_{i\kappa}$ (cf. Lemma 3.4) can be reformulated in terms of the sesquilinear form $a(\cdot,\cdot)$ as
@@
(13) $a(\boldsymbol{u},S_{i\kappa}\boldsymbol{u})\geq C_{A}\left\lVert{\boldsymbol{u}}\right\rVert^{2}_{\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)},$
for all $\boldsymbol{u}\in\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$. The following generalized Gårding inequality for $b(\cdot,\cdot)$ is an immediate consequence of the inequality (13).
@@
For $\kappa>0$, it holds for all $\boldsymbol{u}\in\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$ that
(14) $\left|{b(\boldsymbol{u},S_{i\kappa}\boldsymbol{u})-t(\boldsymbol{u},S_{i\kappa}\boldsymbol{u})}\right|\geq C_{A}\left\lVert{\boldsymbol{u}}\right\rVert^{2}_{\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)},$
with $C_{A}>0$ the constant defined in Lemma 3.4 depending only on $\Gamma$ and $\kappa$.
Finally, the unique solvability of the variational problem (10) follows from the injectivity of the MFIE (i.e., Assumption 1) and the generalized Gårding inequality (14), by means of a Fredholm alternative argument (cf. [5, Proposition 3]).
@@ 5 Galerkin discretization
In this section, we analyze a Galerkin boundary element discretization scheme for the variational formulation (10) that was firstly introduced in [15]. For the sake of convenience, from now on, we use the shorter notation $\operatorname{\mathbf{X}}_{\Gamma}$ for the trace space $\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$, i.e,
$\operatorname{\mathbf{X}}_{\Gamma}:=\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma).$
@@ 5.1 Galerkin boundary element discretization
$\operatorname{\mathbf{U}}_{h}:=\left\{{\boldsymbol{u}_{h}\in\operatorname{\mathbf{X}}_{\Gamma}:\left.\boldsymbol{u}_{h}\right|_{T}\in\operatorname{RT}_{0}(T),\,\forall T\in\mathcal{T}_{h}}\right\},$
@@
for all $e\in\mathcal{E}_{h}$, where $\textbf{n}_{j}$ is the normal of a face $\Gamma_{j}$ in whose the closure of $e$ is contained. The boundary element space $\operatorname{\mathbf{U}}_{h}$ is also called the Rao-Wilton-Glission space [30].
@@
The Raviart-Thomas boundary element spaces have been widely used in the Galerkin discretization of the EFIE and CFIEs, both for the solution space and the test space, see, e.g., [22, 8, 25]. These discretization schemes are primarily based on the ellipticity of the operators $S_{\kappa}$ and $S_{i\kappa}$. Unfortunately, this method fails when applied to the MFIE, even on smooth surfaces. It has been shown in [11, Section 3.1] that there exist $\alpha>0$ and $\operatorname{\mathbf{W}}_{h}\subset\operatorname{\mathbf{U}}_{h}$ such that, for all $\boldsymbol{w}_{h}\in\operatorname{\mathbf{W}}_{h}$
$\sup_{\boldsymbol{u}_{h}\in\operatorname{\mathbf{U}}_{h}\setminus\left\{{\boldsymbol{0}}\right\}}\dfrac{\left|{d(\boldsymbol{w}_{h},\boldsymbol{u}_{h})}\right|}{\left\lVert{\boldsymbol{u}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\Gamma}}}\leq\alpha h^{-1/2}\left\lVert{\boldsymbol{w}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\Gamma}}.$
In other words, the inf-sup condition (11) does not hold true on the discrete subspace $\operatorname{\mathbf{U}}_{h}\subset\operatorname{\mathbf{X}}_{\Gamma}$, or equivalently $\operatorname{\mathbf{U}}_{h}$ is not dual to itself with respect to the pairing ,×$\left\langle{\cdot,\cdot}\right\rangle_{\times,\Gamma}$. In fact, there is ample numerical evidence that for a very broad class of situations that are of practical interest, the discretization matrix of the sesquilinear form $d$ using the boundary element space $\operatorname{\mathbf{U}}_{h}$ is singular, with a nullspace whose dimension is a sizeable fraction of the dimension of $\operatorname{\mathbf{U}}_{h}$ [16].
In order to preserve the inf-sup condition for the Galerkin discretization of the sesquilinear form $d(\cdot,\cdot)$, the authors of [15] proposed the test space $\operatorname{\mathbf{V}}_{h}\subset\operatorname{\mathbf{X}}_{\Gamma}$ consisting of Buffa-Christiansen basis functions, which were introduced in [2] (see Figure 1). More particularly, $\dim\operatorname{\mathbf{V}}_{h}=\dim\operatorname{\mathbf{U}}_{h}$ and the following duality between $\operatorname{\mathbf{U}}_{h}$ and $\operatorname{\mathbf{V}}_{h}$ holds
(15) $\sup_{\boldsymbol{u}_{h}\in\operatorname{\mathbf{U}}_{h}\setminus\left\{{\boldsymbol{0}}\right\}}\dfrac{\left|{d(\boldsymbol{u}_{h},\boldsymbol{v}_{h})}\right|}{\left\lVert{\boldsymbol{u}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\Gamma}}}\geq\widetilde{C}_{D}\left\lVert{\boldsymbol{v}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\Gamma}},$
for all $\boldsymbol{v}_{h}\in\operatorname{\mathbf{V}}_{h}$, for some constant $\widetilde{C}_{D}>0$ independent of $h$. Then, the Galerkin boundary element discretization of the variational formulation (10) reads as: find $\boldsymbol{u}_{h}\in\operatorname{\mathbf{U}}_{h}$ such that, for all $\boldsymbol{v}_{h}\in\operatorname{\mathbf{V}}_{h}$
@@
(16) $b(\boldsymbol{u}_{h},\boldsymbol{v}_{h})=\left\langle{\boldsymbol{f},\overline{\boldsymbol{v}_{h}}}\right\rangle_{\times,\Gamma}.$
@@
Figure 1: A Buffa-Christiansen basis element expressed as a linear combination of Raviart-Thomas basis elements on the barycentric refinement $\Gamma_{h}^{\prime}$ of the triangulation $\Gamma_{h}$, with coefficient of each edge multiplied by the number indicated at its origin. The edges of the refinement $\Gamma_{h}^{\prime}$ are oriented away from the central edge.
The following lemma presents a discrete analogue to (13).
@@
There exist $h_{0}>0$ and a constant $\widetilde{C}_{A}>0$ depending only on $\Gamma$ and $\kappa$ such that, for all $h<h_{0}$, there is an isomorphism $T_{h}:\operatorname{\mathbf{U}}_{h}\to\operatorname{\mathbf{V}}_{h}$ such that, for all $\boldsymbol{u}_{h}\in\operatorname{\mathbf{U}}_{h}$
(17) $a(\boldsymbol{u}_{h},T_{h}\boldsymbol{u}_{h})\geq\widetilde{C}_{A}\left\lVert{\boldsymbol{u}_{h}}\right\rVert^{2}_{\operatorname{\mathbf{X}}_{\Gamma}}.$
@@
Let us define the operators $I_{h}:\operatorname{\mathbf{V}}_{h}\to\operatorname{\mathbf{U}}_{h}^{\prime}$ and $S_{h}:\operatorname{\mathbf{U}}_{h}\to\operatorname{\mathbf{U}}_{h}^{\prime}$ by
$\left\langle{I_{h}\boldsymbol{v}_{h},\boldsymbol{w}_{h}}\right\rangle_{\operatorname{\mathbf{U}}_{h}^{\prime}\times\operatorname{\mathbf{U}}_{h}}$ $:=\left\langle{\boldsymbol{w}_{h},\overline{\boldsymbol{v}_{h}}}\right\rangle_{\times,\Gamma},$
$\left\langle{S_{h}\boldsymbol{u}_{h},\boldsymbol{w}_{h}}\right\rangle_{\operatorname{\mathbf{U}}_{h}^{\prime}\times\operatorname{\mathbf{U}}_{h}}$ $:=\left\langle{\boldsymbol{w}_{h},S_{i\kappa}\overline{\boldsymbol{u}_{h}}}\right\rangle_{\times,\Gamma},$
for all $\boldsymbol{u}_{h},\boldsymbol{w}_{h}\in\operatorname{\mathbf{U}}_{h}$ and all $\boldsymbol{v}_{h}\in\operatorname{\mathbf{V}}_{h}$. Due to the $\operatorname{\mathbf{X}}_{\Gamma}$-ellipticity of $S_{i\kappa}$ (cf. Lemma 3.3) and the duality (15), the operator $I_{h}^{-1}\circ S_{h}$ forms an isomorphism between $\operatorname{\mathbf{U}}_{h}$ and $\operatorname{\mathbf{V}}_{h}$. In addition, it holds for all $\boldsymbol{u}_{h},\boldsymbol{w}_{h}\in\operatorname{\mathbf{U}}_{h}$ that
(18) $\left\langle{\boldsymbol{w}_{h},(I_{h}^{-1}\circ S_{h})\overline{\boldsymbol{u}_{h}}}\right\rangle_{\times,\Gamma}=\left\langle{\boldsymbol{w}_{h},S_{i\kappa}\overline{\boldsymbol{u}_{h}}}\right\rangle_{\times,\Gamma}.$
Based on this relation and the uniform discrete inf-sup condition (15), we can deduce for all $\boldsymbol{u}_{h}\in\operatorname{\mathbf{U}}_{h}$ that
(19) $\widetilde{C}_{D}\left\lVert{(I_{h}^{-1}\circ S_{h})\boldsymbol{u}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\Gamma}}$ $\leq\sup_{\boldsymbol{w}_{h}\in\operatorname{\mathbf{U}}_{h}\setminus\left\{{\boldsymbol{0}}\right\}}\dfrac{\left|{\left\langle{\boldsymbol{w}_{h},(I_{h}^{-1}\circ S_{h})\overline{\boldsymbol{u}}_{h}}\right\rangle_{\times,\Gamma}}\right|}{\left\lVert{\boldsymbol{w}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\Gamma}}}$
@@
$=\sup_{\boldsymbol{w}_{h}\in\operatorname{\mathbf{U}}_{h}\setminus\left\{{\boldsymbol{0}}\right\}}\dfrac{\left|{\left\langle{\boldsymbol{w}_{h},S_{i\kappa}\overline{\boldsymbol{u}_{h}}}\right\rangle_{\times,\Gamma}}\right|}{\left\lVert{\boldsymbol{w}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\Gamma}}}$
$\leq C\left\lVert{\boldsymbol{u}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\Gamma}}.$
Next, for all $\boldsymbol{u}_{h},\boldsymbol{w}_{h}\in\operatorname{\mathbf{U}}_{h}$ and for all $\boldsymbol{w}\in\operatorname{\mathbf{X}}_{\Gamma}$, we invoke the relation (18) again to obtain that
$\left\langle{\boldsymbol{w},(I_{h}^{-1}\circ S_{h})\overline{\boldsymbol{u}_{h}}-S_{i\kappa}\overline{\boldsymbol{u}_{h}}}\right\rangle_{\times,\Gamma}$ $=\left\langle{\boldsymbol{w}-\boldsymbol{w}_{h},(I_{h}^{-1}\circ S_{h})\overline{\boldsymbol{u}_{h}}-S_{i\kappa}\overline{\boldsymbol{u}_{h}}}\right\rangle_{\times,\Gamma}$
$\leq\left\lVert{d}\right\rVert\left\lVert{(I_{h}^{-1}\circ S_{h})\overline{\boldsymbol{u}_{h}}-S_{i\kappa}\overline{\boldsymbol{u}_{h}}}\right\rVert_{\operatorname{\mathbf{X}}_{\Gamma}}\left\lVert{\boldsymbol{w}-\boldsymbol{w}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\Gamma}}$
$\leq\widetilde{C}\left\lVert{\boldsymbol{u}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\Gamma}}\left\lVert{\boldsymbol{w}-\boldsymbol{w}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\Gamma}},$
[…]

### v2

@@ 4 The magnetic field integral equation
where $\boldsymbol{u}:=\gamma_{N}^{+}\boldsymbol{e}\in\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$ is the unknown exterior Neumann trace of the scattered electric field, and $\boldsymbol{f}\in\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$ is given via the Dirichlet boundary condition (4) as
@@
4.1 Flawed idea
One may think of deriving the variational formulation of (13) using the $\operatorname{\mathbf{X}}_{\kappa^{\prime}}$-inner product, i.e.,
@@ 4.1 Flawed idea
(14) $\left\langle{\left({\dfrac{1}{2}Id+C_{\kappa}}\right)\boldsymbol{u},S_{i\kappa^{\prime}}\overline{\boldsymbol{v}}}\right\rangle_{\times,\Gamma}=\left\langle{\boldsymbol{f},S_{i\kappa^{\prime}}\overline{\boldsymbol{v}}}\right\rangle_{\times,\Gamma},\qquad\quad\forall\boldsymbol{v}\in\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma).$
Due to Lemmas 3.2 and 3.7, the governing MFIE operator is compact perturbation of a $\operatorname{\mathbf{X}}_{\kappa^{\prime}}$-elliptic operator, for any $\kappa^{\prime}>0$. Given that it is injective (i.e., Assumption 1 is satisfied), the variational formulation (14) has a unique solution $\boldsymbol{u}\in\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$. Moreover, the Galerkin method converges quasi-optimally when applied to this formulation, see [37, Theorem 4.2.9].
Despite that, the equation (14) involves three boundary integrals rather than two as in standard variational formulations for boundary integral equations (e.g., for EFIE in [26]). Computing the entries of the Galerkin matrices for (14) requires the computation of integrals over all of $\Gamma$, even when the basis functions have support on small subsets of $\Gamma$. This disadvantage leads the assembling time of (14) even exceeding the time for solving it, making the formulation (14) undesired in practice.
4.2 Variational formulation
The appropriate variational formulation of (13) is with respect to the duality pairing ,×$\left\langle{\cdot,\cdot}\right\rangle_{\times,\Gamma}$: find $\boldsymbol{u}\in\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$ such that
@@ 4.2 Variational formulation
(15) $b(\boldsymbol{u},\boldsymbol{v})=\left\langle{\boldsymbol{f},\overline{\boldsymbol{v}}}\right\rangle_{\times,\Gamma},\qquad\qquad\quad\forall\boldsymbol{v}\in\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma),$
where sesquilinear form $b:\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)\times\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)\to\mathbb{C}$ is defined by
$b(\boldsymbol{u},\boldsymbol{v}):=\left\langle{\left({\dfrac{1}{2}Id+C_{\kappa}}\right)\boldsymbol{u},\overline{\boldsymbol{v}}}\right\rangle_{\times,\Gamma},\qquad\qquad\boldsymbol{u},\boldsymbol{v}\in\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma).$
On smooth surfaces, it is well-known that the operator $C_{\kappa}$ is compact, see, e.g., [33, Section 5.4] or [8, Lemma 11]. As a result, the self-duality of the space $\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$ with respect to ,×$\left\langle{\cdot,\cdot}\right\rangle_{\times,\Gamma}$ ensures the coercivity of the sesquilinear form $b(\cdot,\cdot)$. More precisely, for any $\kappa^{\prime}>0$, it holds that
(16) $\sup_{\boldsymbol{v}\in\operatorname{\mathbf{X}}_{\kappa^{\prime}}}\dfrac{\left|{\left\langle{\boldsymbol{u},\overline{\boldsymbol{v}}}\right\rangle_{\times,\Gamma}}\right|}{\left\lVert{S^{-1}_{i\kappa^{\prime}}\boldsymbol{v}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}}=\left\lVert{\boldsymbol{u}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}},\qquad\qquad\forall\boldsymbol{u}\in\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma).$
To prove this identity, one just need to take into account the fact that
$\left|{\left\langle{\boldsymbol{u},\overline{\boldsymbol{v}}}\right\rangle_{\times,\Gamma}}\right|=\left|{\left\langle{\boldsymbol{u},\left({S_{i\kappa^{\prime}}\circ S^{-1}_{i\kappa^{\prime}}}\right)\overline{\boldsymbol{v}}}\right\rangle_{\times,\Gamma}}\right|\leq\left\lVert{\boldsymbol{u}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}\left\lVert{S^{-1}_{i\kappa^{\prime}}\boldsymbol{v}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}},$
for all $\boldsymbol{u},\boldsymbol{v}\in\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$, and that the equality occurs when $\boldsymbol{v}=S_{i\kappa^{\prime}}\boldsymbol{u}$. The relation (16) implies that both the inf-sup and the boundedness constants (with respect to the norms specified in (16)) of the duality pairing ,×$\left\langle{\cdot,\cdot}\right\rangle_{\times,\Gamma}$ are 1.
On polyhedral boundaries, the compactness of $C_{\kappa}$ does not necessarily hold true. In this paper, we exploit the compactness of the operator $C_{\kappa}-C_{i\kappa^{\prime}}$ (cf. Lemma 3.2) and the $S_{i\kappa^{\prime}}$-coercivity stated in (12) to prove the coercivity of the sesquilinear form $b(\cdot,\cdot)$. For that reason, we split $b(\cdot,\cdot)$ into a principal part $a(\cdot,\cdot)$ and a compact sesquilinear form $t(\cdot,\cdot)$ as follows
$a(\boldsymbol{u},\boldsymbol{v}):=\left\langle{\left({\dfrac{1}{2}Id+C_{i\kappa^{\prime}}}\right)\boldsymbol{u},\overline{\boldsymbol{v}}}\right\rangle_{\times,\Gamma},\qquad\quad t(\boldsymbol{u},\boldsymbol{v}):=\left\langle{\left({C_{\kappa}-C_{i\kappa^{\prime}}}\right)\boldsymbol{u},\overline{\boldsymbol{v}}}\right\rangle_{\times,\Gamma},$
for some $\kappa^{\prime}>0$. The $S_{i\kappa^{\prime}}$-coercivity of the operator $\tfrac{1}{2}Id+C_{i\kappa^{\prime}}$ (cf. Lemma 3.7) can be reformulated in terms of the sesquilinear form $a(\cdot,\cdot)$ as
(17) $a(\boldsymbol{u},S_{i\kappa^{\prime}}\boldsymbol{u})\geq\left({\dfrac{1}{4}-\beta^{2}}\right)\left\lVert{\boldsymbol{u}}\right\rVert^{2}_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}},\qquad\qquad\forall\boldsymbol{u}\in\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma).$
The following generalized Gårding inequality for $b(\cdot,\cdot)$ is an immediate consequence of the inequality (17).
@@
For any $\kappa,\kappa^{\prime}>0$, it holds that
(18) $\left|{b(\boldsymbol{u},S_{i\kappa^{\prime}}\boldsymbol{u})-t(\boldsymbol{u},S_{i\kappa^{\prime}}\boldsymbol{u})}\right|\geq\left({\dfrac{1}{4}-\beta^{2}}\right)\left\lVert{\boldsymbol{u}}\right\rVert^{2}_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}},\qquad\forall\boldsymbol{u}\in\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma).$
Finally, the unique solvability of the variational problem (15) follows from the injectivity of the MFIE (i.e., Assumption 1) and the generalized Gårding inequality (18), by means of a Fredholm alternative argument (cf. [6, Proposition 3]).
@@ 5 Galerkin discretization
In this section, we analyze the Galerkin boundary element discretization scheme for the variational formulation (15) that was firstly introduced in [17].
@@ 5.1 Galerkin boundary element discretization
$\operatorname{\mathbf{U}}_{h}:=\left\{{\boldsymbol{u}_{h}\in\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma):\left.\boldsymbol{u}_{h}\right|_{T}\in\operatorname{RT}_{0}(T),\,\forall T\in\mathcal{T}_{h}}\right\},$
@@
for all $e\in\mathcal{E}_{h}$, where $\textbf{n}_{j}$ is the normal of a face $\Gamma_{j}$ in whose the closure of $e$ is contained. The boundary element space $\operatorname{\mathbf{U}}_{h}$ is also called the Rao-Wilton-Glisson space [34].
@@
The Raviart-Thomas boundary elements have been widely used in the Galerkin discretization of the EFIE and CFIEs, both for the solution space and the test space, see, e.g., [26, 9]. These discretization schemes are primarily based on the ellipticity of the operator $S_{\kappa}$. Unfortunately, this method fails when applied to the MFIE, even on smooth surfaces. It has been shown in [13, Section 3.1] that there exist $C>0$ and $\operatorname{\mathbf{W}}_{h}\subset\operatorname{\mathbf{U}}_{h}$ such that
@@
$\sup_{\boldsymbol{u}_{h}\in\operatorname{\mathbf{U}}_{h}}\sup_{\boldsymbol{w}_{h}\in\operatorname{\mathbf{W}}_{h}}\dfrac{\left|{\left\langle{\boldsymbol{u}_{h},\overline{\boldsymbol{w}_{h}}}\right\rangle_{\times,\Gamma}}\right|}{\left\lVert{\boldsymbol{u}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}\left\lVert{S^{-1}_{i\kappa^{\prime}}\boldsymbol{w}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}}\leq Ch^{-1/2}.$
In other words, the subspace $\operatorname{\mathbf{U}}_{h}$ is not dual to itself with respect to the pairing ,×$\left\langle{\cdot,\cdot}\right\rangle_{\times,\Gamma}$. In fact, there is ample numerical evidence that for a very broad class of situations that are of practical interest, the discretization matrix of the anti-symmetric pairing ,×$\left\langle{\cdot,\cdot}\right\rangle_{\times,\Gamma}$ using the boundary element space $\operatorname{\mathbf{U}}_{h}$ is singular, with a nullspace whose dimension is a sizeable fraction of the dimension of $\operatorname{\mathbf{U}}_{h}$ [18].
In order to preserve the inf-sup condition for the Galerkin discretization of ,×$\left\langle{\cdot,\cdot}\right\rangle_{\times,\Gamma}$, the authors of [17] proposed the test space $\operatorname{\mathbf{V}}_{h}\subset\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$ consisting of the Buffa-Christiansen boundary elements [3] (see Figure 1). More particularly, $\dim\operatorname{\mathbf{V}}_{h}=\dim\operatorname{\mathbf{U}}_{h}$ and the following duality between $\operatorname{\mathbf{U}}_{h}$ and $\operatorname{\mathbf{V}}_{h}$ holds
(19) $\inf_{\boldsymbol{u}_{h}\in\operatorname{\mathbf{U}}_{h}}\sup_{\boldsymbol{v}_{h}\in\operatorname{\mathbf{V}}_{h}}\dfrac{\left|{\left\langle{\boldsymbol{u}_{h},\overline{\boldsymbol{v}_{h}}}\right\rangle_{\times,\Gamma}}\right|}{\left\lVert{\boldsymbol{u}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}\left\lVert{S^{-1}_{i\kappa^{\prime}}\boldsymbol{v}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}}\geq\alpha,$
for some constant $\alpha(\kappa^{\prime},\Gamma)>0$ independent of $h$, with $\kappa^{\prime}>0$. Obviously, $\alpha\leq 1$ due to (16). In what follows, $\alpha$ represents the maximal value that still fulfills (19). Then, the Petrov-Galerkin boundary element discretization of the variational formulation (15) reads as: find $\boldsymbol{u}_{h}\in\operatorname{\mathbf{U}}_{h}$ such that
(20) $b(\boldsymbol{u}_{h},\boldsymbol{v}_{h})=\left\langle{\boldsymbol{f},\overline{\boldsymbol{v}_{h}}}\right\rangle_{\times,\Gamma},\qquad\qquad\quad\forall\boldsymbol{v}_{h}\in\operatorname{\mathbf{V}}_{h}.$
@@
Figure 1: A Buffa-Christiansen basis element expressed as a linear combination of Raviart-Thomas basis elements on the barycentric refinement $\widetilde{\Gamma}_{h}$ of the triangulation $\Gamma_{h}$, with coefficient of each edge multiplied by the number indicated at its origin. The edges of the refinement $\widetilde{\Gamma}_{h}$ are oriented away from the central edge.
The following lemma presents the discrete inf-sup condition for the sesquilinear form $a(\cdot,\cdot)$ restricted on $\operatorname{\mathbf{U}}_{h}\times\operatorname{\mathbf{V}}_{h}$. This condition is a weaker discrete counterpart of the $S_{i\kappa^{\prime}}$-coercivity (17).
@@
For any $\kappa^{\prime}>0$, the following inf-sup condition is satisfied
(21) $\inf_{\boldsymbol{u}_{h}\in\operatorname{\mathbf{U}}_{h}}\sup_{\boldsymbol{v}_{h}\in\operatorname{\mathbf{V}}_{h}}\dfrac{\left|{a(\boldsymbol{u}_{h},\boldsymbol{v}_{h})}\right|}{\left\lVert{\boldsymbol{u}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}\left\lVert{S^{-1}_{i\kappa^{\prime}}\boldsymbol{v}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}}\geq\dfrac{\alpha}{2}-\beta.$
@@
The following estimate holds true due to the discrete inf-sup condition (19) and the contraction property (9)
@@
$\inf_{\boldsymbol{u}_{h}\in\operatorname{\mathbf{U}}_{h}}\sup_{\boldsymbol{v}_{h}\in\operatorname{\mathbf{V}}_{h}}\dfrac{\left|{a(\boldsymbol{u}_{h},\boldsymbol{v}_{h})}\right|}{\left\lVert{\boldsymbol{u}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}\left\lVert{S^{-1}_{i\kappa^{\prime}}\boldsymbol{v}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}}$ $\geq\dfrac{1}{2}\inf_{\boldsymbol{u}_{h}\in\operatorname{\mathbf{U}}_{h}}\sup_{\boldsymbol{v}_{h}\in\operatorname{\mathbf{V}}_{h}}\dfrac{\left|{\left\langle{\boldsymbol{u}_{h},\overline{\boldsymbol{v}}_{h}}\right\rangle_{\times,\Gamma}}\right|}{\left\lVert{\boldsymbol{u}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}\left\lVert{S^{-1}_{i\kappa^{\prime}}\boldsymbol{v}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}}$